6. The expected number of typographical errors on a page of a certain magazine is H O50. What is the probability that the next page you read contains: 0167T a) exactly one typographical error? b) at least one typographical error?
Answer #1

Let X denotes the number of typographical errors on a randomly selected page.

X ~ Poisson(mu = 0.75)

The probability mass function of X is

0.60.67

a) The probability that there is exactly one typographical error

P(X = 1)

一0.67 * 0.671

= 0.342845

b) The probability that there is at least one typographical error

=P(Xgeq 1)

= 1 - P(X = 0)

= 1 - 0.511709

= 0.488291